Grade 10Physical Chemistry

Two small stones are tied at the ends of a rope of length 5 m. Imagine you are
holding them, one in each hand, with the rope hanging loosely on the top floor
balcony of a tall building. You release one stone at a certain moment and the
other 0.5 s later along the same vertical line from the same position. The rope
will be taut after a time of __________

jui
11 Years agoGrade 10

1 Answer

Abhishek Singh
11 Years ago
for first stone writing eqn of motion 
x1= ½gt2
for second stone, 
x2=1/2g(t-0.5)2   for t.0.5s
string becomes taut at the instant 
x1-x2=5
½ gt2-1/2g(t-0.5)2=5
½g(0.5)(2t-0.5)=5
t=5/4 seconds